OpenStudy (anonymous):

If sin Θ = negative square root 3 over 2 and π < Θ < 3 pi over 2, what are the values of cos Θ and tan Θ?

OpenStudy (gamer56):

Do you know how to simplify?

OpenStudy (anonymous):

This lesson doesn't really make any sense to me.

OpenStudy (gamer56):

Does this help you any? Its an example~ :D (sinx - cosx)^2 - (sinx + cosx)^2 = (sin^2 x - 2sinxcosx + cos^2 x) - (sin^2 x + 2sinxcosx + cos^2 x) = - 2 sinx cosx - 2 sinx cosx = - 4 sinx cosx = - 2sin(2x) sin(2pi/3) = sqrt(3)/2

OpenStudy (anonymous):

Eh, a little bit. I have the rations for sin, cos, and tan, idk if I would use it though

OpenStudy (anonymous):

ratios*

OpenStudy (gamer56):

Yes you should use. Oh and if I helped you medal and fan please?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Anybody else can help/guide me through this question?

OpenStudy (anonymous):

I think it's A, but idk, anyone?

OpenStudy (anonymous):

@gamer56, could it be a?

OpenStudy (gamer56):

First of all,I cant see your answer choices

OpenStudy (anonymous):

O, lol

OpenStudy (anonymous):

cos Θ = negative 1 over 2; tan Θ = square root 3 cos Θ = negative 1 over 2; tan Θ = −1 cos Θ = square root 3 over 4; tan Θ = −2 cos Θ = 1 over 2; tan Θ = square root 3

OpenStudy (gamer56):

:D

OpenStudy (anonymous):

So, is it A or...

OpenStudy (anonymous):

Alright, it is A, thank you @gamer56!

OpenStudy (gamer56):

@flashy32 welcome fan and medal please?